OpenStudy (anonymous):

John is 3 times as old as his son, Michael, and in 'x' years' time, he will be twice and old as him. (i) Write down expressions for John's and Michael's age in 'x' years' time. (ii) Write down, and solve, an equation in 'x'.

OpenStudy (anonymous):

let John's present age be 'j' and Michael's present age be 'm' j=3m-------(i) after x years John=j+x Michael=m+x j+x=2(m+x)------(ii) substituting (i) in (ii) 3m+x=2(m+x) 3m+x=2m+2x 3m-2m=2x-x m=x